Hi, can you help me by solving the equation: 5[2x-5]+3=18
(The ''['' are absolute value)
Thank you.

Answer by stanbon(75887)

5[2x-5]+3=18
----------
|2x-5| = 3
------
2x-5 = 3 or 2x-5 = -3
-----------------------------
x = 4 or x = 1
-------------------
Cheers,
Stan H.
---------------